# Explicit vs. Implicit: Investigating Social Bias in Large Language Models through Self-Reflection

### v2

@@ 4.1 Experimental Setup
This study systematically examined six types of typical social stereotypes to comprehensively assess bias in LLMs. Four categories are derived from classical IAT Greenwald et al. (1998) and WEAT Caliskan et al. (2017) studies, utilizing their validated target words, attribute words, and corresponding stimuli. These four categories include:
1. Gender stereotypes in career-family domains: The strength of association between male/female and career/family.
2. Racial stereotypes: The degree of association between White/Black people and positive/negative emotions.
@@
4. Gender stereotypes in academic domains: The degree of association between male/female and science/arts.
Beyond these four categories, this study added two important research subjects:
5. Disability stereotypes: The association between able-bodied/disabled and positive/negative emotions. Data for this part of research is based on the official Implicit Association Test website [^1] https://implicit.harvard.edu/implicit/takeatest.html
6. Occupational gender stereotypes: The degree of association between male/female and male-dominated/female-dominated occupations . Given the prevalence and research value of occupational gender bias studies in natural language processing Smith et al. (2022); Watson et al. (2023), this study collected 10 pairs of semantically similar occupations with significant gender distribution differences from the U.S. Bureau of Labor Statistics website [^2] https://www.bls.gov/cps/cpsaat11.htm. Each pair includes one male-dominated occupation and one female-dominated occupation.
For both implicit and explicit bias tests, we conducted 20 independent experiments on each of the 10 test templates per research subject, totaling 200 experiments for each type of bias. Given that this study involved 9 large language models, 6 types of stereotypes, and 2 types of bias (implicit and explicit), the total number of experiments reached 21,600 (9 × 6 × 200 × 2), a scale sufficient to ensure statistical reliability of the results. Following the research findings of Wang et al. Wang et al. (2023), all model evaluation parameters are set with Temperature=0 to reduce output randomness and obtain more reliable assessment results.
@@ 4.2 Experimental Metrics
Stereotypes refer to fixed, simplified, and often biased views or beliefs about a particular group or individualDovidio (2010). These views are typically based on characteristics such as race, gender, age, and occupation while often disregarding individual differences Gilbert and Hixon (1991). Our study focuses on examining the stereotypes reflected in both explicit and implicit biases in LLMs, and defines corresponding quantitative metrics.
In measuring implicit bias in this study, a stereotype is considered present when the LLM establishes semantic associations between each element of a pair of attributes and its corresponding traditional stereotypical target; otherwise, it is considered absent. This strict definition criterion helps identify the LLM’s implicit biases. For example, when the LLM associates White stimuli (such as White names) with positive attributes and Black stimuli (such as Black names) with negative attributes, it is considered to exhibit stereotyping.
In measuring explicit bias, as this study employs a five-point Likert scaleLikert (1932) for evaluation, stereotype expression statements are constructed based on the implicit bias template sentences, associating each element of a pair of attributes with its traditional stereotypical target, and requiring the LLM to make judgments. When the model selects "agree" or "strongly agree" and provides corresponding reasons, it is considered to exhibit stereotyping. When the LLM selects "neutral," "disagree," or "strongly disagree," it is considered to show no stereotyping.
To quantify the experimental results, this paper defines the "Stereotypical Score" (SC):
$Stereotypical\ Score=\frac{n_{stereotype}}{N}$ (1)
where $n_{stereotype}$ represents the number of stereotype expressions occurring in either explicit or implicit tests, and $N$ is the total number of experiments. This metric ranges from $[0,1]$, with scores closer to $1$ indicating more severe stereotyping and scores closer to 00 indicating less stereotyping.
@@ 4.3 Baselines
This study first selects high-ranking large language models from the Chatbot Arena platform Chiang et al. (2024) as research subjects. Chatbot Arena is an open language model evaluation platform that employs human preference-based assessment through a paired comparison mechanism: when users pose questions, the platform randomly selects two anonymous language models to respond, and users vote based on the quality of the models’ responses, with model identities revealed only after voting is complete. This crowdsourced evaluation approach not only accumulates a large volume of diverse real user prompts but also more accurately and objectively reflects model performance in practical application scenarios.
Specifically, among mainstream commercial models, this paper selected the following representative models: GPT-4-turboHurst et al. (2024), GPT-4oHurst et al. (2024), Claude-3.5-SonnetBai et al. (2022b), Gemini-2.0-FlashTeam et al. (2023), and others. To enhance the comprehensiveness of this experiment, this paper also included LLMs with significant influence in the open-source community: LLaMA-2-70B-chatTouvron et al. (2023) and LLaMA-3.1-405B-InstructDubey et al. (2024). Table 1 presents detailed information about these selected LLMs.
@@
Table 1: Comparison of Different prominent Large Language Models
To further investigate the factors influencing explicit and implicit biases of LLMs, this study selects the LLaMA series of open-source LLMs as experimental subjects. These LLMs not only provide transparent information regarding their pre-training data scale and parameters but also offer both pre-trained base models and instruction-tuned versions, providing ideal conditions for studying the independent effects and interactions of different factors.
In examining the impact of training data scale and model parameters on both types of bias, this paper selected the pre-trained versions of LLaMA-2, LLaMA-3, LLaMA-3.1, and LLaMA-3.2 series models. These LLMs range in parameter size from 1B to 405B, with training data spanning from 1T tokens to over 15T tokens, providing extensive experimental comparison conditions. The detailed information about these selected LLMs is shown in Table 2
@@
Table 2: Configurations of LLaMA model series.
In studying the effects of alignment training on explicit and implicit biases, this research focused primarily on the 1B and 3B models from the LLaMA-3.2 series. These LLMs, with their smaller parameter, have relatively lower computational resource requirements, making them more suitable for multiple rounds of preference alignment training experiments.
Specifically, this study employs Direct Preference Optimization (DPO)Rafailov et al. (2024) combined with Low-Rank Adaptation (LoRA)Hu et al. (2022), conducting human preference alignment training at various levels on Anthropic’s open-source harmless-base RLHF datasetBai et al. (2022a). Compared to Reinforcement Learning from Human Feedback (RLHF)Ouyang et al. (2022), Direct Preference Optimization (DPO) offers a more streamlined training process by eliminating the need for complex reward models and reinforcement learning procedures, resulting in significantly simplified implementation and reduced computational overhead. To comprehensively evaluate the impact of preference alignment training on model behavior, this study implements a fine-grained training step design, conducting evaluations at 100-step intervals until observing the near-elimination of one type of bias. The alignment training is performed on a single NVIDIA RTX 4090 GPU. To balance computational resources and training effectiveness, a gradient accumulation strategy is adopted: setting the batch size to 32, processing 4 samples per iteration with 8 gradient accumulation steps. The temperature coefficient $\beta$ in DPO is set to 0.1.
@@ 5 Results and Analysis
Figure 2: Average Stereotype Scores: Comparing Explicit and Implicit Bias Across LLMs.
Table 3 presents detailed stereotype scores for explicit and implicit biases across six dimensions: age, disability, gender career, gender occupation, race, and science. Figure 2 visualizes the overall explicit and implicit bias performance across LLMs, ordered by model capability based on Chatbot Arena platform. The data reveals a consistent pattern across various LLMs: a notable inconsistency between implicit and explicit biases across different attribute dimensions. While implicit biases consistently show strong stereotypical tendencies, explicit biases remain relatively mild. This inconsistency is evident across multiple attributes. Furthermore, analysis of the relation between model capability and bias performance reveals a significant finding: model general capability positively correlates with the degree of implicit bias. Specifically, models ranking higher on the Chatbot Arena platform tend to exhibit stronger implicit stereotypical biases.
@@ 5.2.1 Impact of Data Count and Parameters
The impact of training data size and model parameters on explicit and implicit biases are illustrated in Figure 3. The experimental results reveal distinctly different trends in explicit and implicit biases as model parameters and training data increase: explicit bias shows a significant decrease with increased training data and parameters, particularly pronounced in large-scale LLM (above 70B). When the parameter scale reaches 405B, explicit bias nearly vanishes (stereotype score of 0.03). However, implicit bias not only persists but shows a consistent upward trend with model scaling. In models with 405B parameters, implicit bias reaches a substantial stereotype score of approximately 0.72.
These findings reveal the persistent nature of implicit bias. Traditional optimization approaches of increasing model parameters and training data size, while effective in mitigating explicit bias, actually intensify implicit bias while improving model performance. This suggests that addressing implicit bias, as a more profound form of bias, may require more targeted solutions rather than merely relying on scaling model size or training data.
@@ 5.2.2 Impact of Alignment Steps
The impact of alignment training on explicit and implicit biases in LLaMA-3.2-1B and LLaMA-3.2-3B are illustrated in Figures 4 and 5.
The experimental results reveal significant differences in how alignment training affects explicit and implicit biases. For explicit bias, alignment training demonstrates a clear inhibitory effect: after 300 training steps, explicit bias shows a rapid decline, eventually approaching zero. This phenomenon is verified in both LLMs, indicating that human preference alignment training effectively reduces explicit bias in LLMs. However, for implicit bias, the effectiveness of alignment training is notably limited. Implicit bias levels remain relatively stable throughout the training process for both 1B and 3B models, maintaining around 0.2. This "resistance to intervention" suggests that traditional human preference alignment methods may struggle to address and improve implicit bias in models. Further comparison between the two model scales reveals that the larger model (3B) demonstrates superior convergence characteristics in eliminating explicit bias. Its explicit bias curve shows significant improvement around 200 steps, while the 1B model requires a longer training period to achieve similar effects.
These findings provide crucial insights into the formation mechanisms of model bias: while alignment training effectively suppresses explicit bias, its intervention on implicit bias remains notably limited. This significant disparity not only reveals the deep complexity of model bias but also indicates the need for future research to develop more targeted approaches in addressing implicit bias.
@@ 6 Conclusion
This paper presents a systematic investigation of explicit and implicit biases in large language models(LLMs), conducting an in-depth comparative analysis of these two forms of bias. Initially, drawing from social psychology theories and methodologies, we propose a self-reflection-based approach to measure both explicit and implicit biases in LLMs. This method establishes explicit bias measurement through models’ self-reflection on implicit bias, enabling comparative analysis of both bias types within the same social groups. Experimental results across multiple LLMs and topics reveal significant inconsistencies between explicit and implicit biases: while LLMs demonstrate low stereotypical tendencies in explicit bias, they exhibit strong stereotypical patterns in implicit bias. Further investigation explores key factors influencing these biases, including training data size, model parameter count, and alignment training. The findings indicate that while increased training data and model parameters significantly reduce explicit bias, implicit bias shows an intensifying trend. Additionally, alignment training (such as RLHF and DPO), while effective in suppressing explicit bias, has limited impact on implicit bias, which maintains relatively stable intensity throughout the alignment process. These systematic experiments reveal the mechanisms behind the divergence between explicit and implicit biases in models, while demonstrating the notable persistence of implicit bias and the limitations of traditional bias mitigation methods in addressing implicit bias issues.

### v4

@@ 4.1 Experimental Setup
This study systematically examines six types of typical social stereotypes to comprehensively assess bias in LLMs. Four categories are derived from classical Implicit Association Test (IAT) Greenwald et al. (1998) and Word Embedding Association Test (WEAT) Caliskan et al. (2017) studies, utilizing their validated target words, attribute words, and corresponding stimuli. These four categories are as follows:
1. Gender stereotypes about career-family: the strength of association between male/female and career/family.
2. Race stereotypes: the degree of association between White/Black people and positive/negative emotions.
@@
4. Gender stereotypes in academic domains: the degree of association between male/female and science/arts disciplines.
Beyond these four categories, this study adds two important research subjects:
5. Disability stereotypes: the association between able-bodied/disabled individuals and positive/negative emotions. Data for this part of research is based on the official IAT website [^1] https://implicit.harvard.edu/implicit/takeatest.html.
6. Occupational gender stereotypes: the degree of association between male/female and male-dominated/female-dominated occupations . Given the prevalence and research value of occupational gender bias studies in NLP Smith et al. (2022); Watson et al. (2023), we collect 10 pairs of semantically similar occupations with significant gender distribution differences from the U.S. Bureau of Labor Statistics website [^2] https://www.bls.gov/cps/cpsaat11.htm. Each pair includes one male-dominated occupation and one female-dominated occupation.
For both implicit and explicit bias tests, we conduct 20 independent experiments on each of the 10 test templates per research subject, totaling 200 experiments for each stereotype category. Given that this study involves 6 LLMs, 6 types of stereotypes, and 2 types of bias (implicit and explicit), the total number of experiments reaches 14,400 (6 × 6 × 200 × 2), a scale sufficient to ensure statistical reliability of experimental results. Following the research findings of Wang et al. (2023), all model evaluation parameters are set with temperature = 0 to reduce output randomness and obtain more reliable assessment results.
@@ 4.2 Experimental Metrics
Stereotypes refer to fixed, simplified, and often biased views or beliefs about a particular individual or group Dovidio (2010). These views are typically based on characteristics such as race, gender, age, and occupation while often disregarding individual differences Gilbert and Hixon (1991). Our study focuses on evaluating the stereotypes reflected in both explicit and implicit biases in LLMs, and defines corresponding quantitative metrics.
In measuring implicit bias, a stereotype is considered present when the LLM establishes semantic associations between each element of a pair of attributes and its corresponding traditional stereotypical target; otherwise, it is considered absent. This strict definition criterion helps identify the LLM’s implicit bias. For example, when the LLM associates stimuli related to White people (such as White names) with positive attributes and stimuli related to Black people (such as Black names) with negative attributes, it is considered to exhibit stereotyping.
In measuring explicit bias, as this study employs a 5-point Likert scale Likert (1932) for evaluation, stereotype expression statements are constructed based on the implicit bias template sentences, associating each element of a pair of attributes with its traditional stereotypical target, and requiring the LLM to make judgments. When the LLM selects "agree" or "strongly agree" and provides corresponding reasons, it is considered to exhibit stereotyping. When the LLM selects "neutral", "disagree", or "strongly disagree", it is considered to show no stereotyping.
To quantify the experimental results, we define the Stereotypical Score (SC):
$SC=\frac{n_{stereotype}}{N},$ (1)
where $n_{stereotype}$ represents the number of stereotypical expressions occurring in either explicit or implicit tests, and $N$ is the total number of experiments. This metric ranges from $[0,1]$, with scores closer to $1$ indicating more severe stereotyping and scores closer to 00 indicating less stereotyping.
@@ 4.3 Baselines
This study first selects high-ranking LLMs from the Chatbot Arena platform Chiang et al. (2024) as research subjects. Chatbot Arena is an open LLM evaluation platform that employs human preference-based assessment through a paired comparison mechanism: when users propose questions, the platform randomly selects two LLMs to respond, and users vote based on the quality of the LLMs’ responses. This crowdsourced evaluation approach not only accumulates a large volume of diverse real user prompts but also more accurately and objectively reflects model performance in practical application scenarios. For our research, we first select several representative mainstream commercial LLMs: GPT-4-Turbo Hurst et al. (2024), GPT-4o Hurst et al. (2024), Claude-3.5-Sonnet Bai et al. (2022b) and Gemini-2.0-Flash Team et al. (2023). To enhance the comprehensiveness of our experiments, we also include LLMs with significant influence in the open-source community: LLaMA-2-70B-chat Touvron et al. (2023) and LLaMA-3.1-405B-Instruct Dubey et al. (2024). Table 1 presents detailed information about these selected LLMs.
@@
Table 1: Comparison of different prominent Large Language Models.
To further investigate the factors influencing explicit and implicit biases of LLMs, we select the LLaMA series of open-source models as experimental subjects. These LLMs not only provide transparent information about pre-training data scale and parameters but also offer both pre-trained base and instruction-tuned versions, providing ideal conditions for studying the independent effects and interactions of different factors.
In examining the impact of training data scale and model parameters on both types of bias, we analyze the pre-trained base versions of LLaMA-2, LLaMA-3, LLaMA-3.1, and LLaMA-3.2 series models. These LLMs range in parameter sizes from 1B to 405B, with training data spanning from 1T tokens to over 15T tokens, providing extensive experimental comparison conditions. The detailed information about these selected LLMs is shown in Table 2.
@@
Table 2: Configurations of LLaMA series.
In studying the effects of alignment training on explicit and implicit biases, we focus primarily on the 1B and 3B pre-trained models from the LLaMA-3.2 series. These LLMs with their smaller parameter sizes require relatively lower computational resources, which makes them more suitable for conducting multiple rounds of preference alignment training experiments. Additionally, we employ Direct Preference Optimization (DPO) Rafailov et al. (2024) combined with Low-Rank Adaptation (LoRA) Hu et al. (2022), conducting human preference alignment training on Anthropic’s open-source harmless-base RLHF dataset Bai et al. (2022a). Compared to Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022), DPO offers a more streamlined training process by eliminating the need for complex reward modeling, resulting in significantly simplified implementation and reduced computational overhead. To comprehensively evaluate the impact of preference alignment training on model behavior, we implement a fine-grained training step design with evaluations at 100-step intervals until observing the near-elimination of one type of bias. All alignment training runs on a single NVIDIA RTX 4090 GPU. We balance computational resources and training effectiveness through gradient accumulation, setting the batch size to 32 by processing 4 samples per iteration with 8 accumulation steps. The temperature coefficient $\beta$ in DPO is set to 0.1.
@@ 5 Results and Analysis
Figure 2: Average stereotypical scores: comparing explicit and implicit biases across different LLMs.
Table 3 presents detailed stereotypical scores for explicit and implicit biases across six dimensions: age, disability, gender career, gender occupation, race, and science. Figure 2 illustrates the average explicit and implicit bias performance across various LLMs, ordered according to their capabilities based on Chatbot Arena platform. These experimental results reveal a consistent pattern across all evaluated LLMs: a notable inconsistency between implicit and explicit biases across different attribute dimensions. While implicit bias shows strong stereotypical tendencies, explicit bias exhibits relatively mild stereotyping. This pattern of inconsistency is evident across multiple attribute categories examined. Furthermore, our analysis of the relation between model capability and bias performance uncovers a significant finding: LLMs’ general capability positively correlates with the degree of implicit bias. Specifically, LLMs ranking higher on the Chatbot Arena platform, indicating superior comprehensive performance, tend to display stronger implicit stereotypes and biases.
@@ 5.2.1 Impact of Data Count and Parameters
Figure 3 illustrates the impact of training data size and model parameters on explicit and implicit biases. The experimental results demonstrate distinctly different trends in explicit and implicit biases as model parameters and training data increase: explicit bias decreases significantly with increased training data and parameters, with this effect being particularly pronounced in large-scale LLM (above 70B). When the parameter scale reaches 405B, explicit bias nearly vanishes (stereotypical score of 0.03). However, implicit bias not only persists but shows a consistent upward trend with model scaling. In models with 405B parameters, implicit bias reaches a substantial stereotypical score of approximately 0.72.
These findings highlight the persistent nature of implicit bias. Traditional optimization approaches of increasing model parameters and training data size, while effective in mitigating explicit bias, actually intensify implicit bias while improving model performance. This suggests that addressing implicit bias, which represents a more profound form of bias, necessitates more targeted solutions rather than merely relying on scaling model size or training data.
@@ 5.2.2 Impact of Alignment Steps
The impact of alignment training on explicit and implicit biases in LLaMA-3.2-1B and LLaMA-3.2-3B is illustrated in Figures 4 and 5.
The experimental results reveal significant differences in how alignment training affects explicit and implicit biases within LLMs. For explicit bias, alignment training demonstrates a clear inhibitory effect: after 300 training steps, explicit bias shows a rapid decline, with its stereotypical score eventually approaching zero. This phenomenon is verified across both LLMs, indicating that human preference alignment training effectively reduces explicit bias. However, for implicit bias, the effectiveness of alignment training is notably limited. Implicit bias levels remain relatively stable throughout the training process for both 1B and 3B models, consistently maintaining around 0.2. This resistance to intervention suggests that conventional human preference alignment methods are ineffective at reducing and addressing implicit bias in models. Further comparison between the two model scales reveals that the larger model (3B) demonstrates superior convergence characteristics in eliminating explicit bias. Its explicit bias curve shows significant decrease around 200 steps, while the 1B model requires a longer training period to achieve comparable effects.
These findings provide crucial insights into the formation mechanisms of model bias: while alignment training effectively suppresses explicit bias, its impact on implicit bias remains notably limited. This significant disparity highlights the deep complexity of model bias and indicates the need for future research to develop more targeted approaches in addressing implicit bias.
@@ 6 Conclusion
In this paper, we present a systematic investigation of explicit and implicit biases within LLMs, conducting an in-depth comparative analysis of these two forms of bias. Drawing from social psychology theories and methodologies, we propose a novel self-reflection-based approach to measure both explicit and implicit biases in LLMs. This method establishes explicit bias measurement through models’ self-reflection on implicit bias, enabling a comparison of both bias types within the same social groups. Our experimental results across multiple LLMs and topics reveal significant inconsistencies between explicit and implicit biases: while LLMs demonstrate low stereotypical tendencies in explicit bias, they exhibit strong stereotypical patterns in implicit bias. We further explore key factors influencing these biases, including training data size, model parameter count, and alignment training. Our findings indicate that while increased training data and model parameters significantly reduce explicit bias, implicit bias shows an intensifying trend. Moreover, while alignment training effectively suppresses explicit bias, its impact on implicit bias remains limited. Through these systematic experiments, we uncover the mechanisms behind the inconsistency between explicit and implicit biases within LLMs. Our results highlight the notable persistence of implicit bias and the limitations of traditional bias mitigation methods in addressing implicit bias issues.
Limitations
First, this paper explores both explicit and implicit biases in LLMs, examining major socially sensitive topics including gender, race, and age. However, these works focus on single-category biases. In reality, biases may exist in intersectional and compound forms. For example, Black female scientists represent a group that involves both racial and gender attributes, and biases toward this group may be different from purely racial bias or gender bias alone. Additionally, when analyzing the inconsistency between explicit and implicit biases, this paper mainly explores three factors: training data size, model parameter count, and alignment training. While these factors indeed have significant impact on model bias, there are other factors that warrant further analysis. For instance, beyond the scale of training data, the sources and types of data also influence model bias; different pre-training tasks or model architectures could also lead to varying degrees of bias. Future research should explore these directions to obtain more complete findings.
Ethical Considerations
This work involves sensitive social dimensions such as race, gender, and disability. We are committed to treating each category with respect and caution, avoiding harmful or offensive language, and ensuring these dimensions are used solely for analyzing model biases. Additionally, the stimuli and prompt materials used in our experiments are sourced from public resources and do not contain personally identifiable information.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (62376188, 62272340, 62276187, 62376192, 62166022) and the Key Technology Research and Industrial Application Demonstration of General Large Model with Autonomous Intelligent Computing Power, No.24ZGZNGX00020.
Appendix A Prompt Templates
The following are the 5 test templates used in our study:
  1. 1.
    <mask> is often to attrX as <mask> is often to attrY.
  2. 2.
    <mask> : attrX, <mask> : attrY.
  3. 3.
    <mask> relates to attrX as <mask> relates to attrY.
  4. 4.
    <mask> is associated with attrX, and <mask> is associated with attrY.
  5. 5.
    <mask> is connected to attrX while <mask> is connected to attrY.
For each template above, we swap the positions of attrX and attrY to mitigate order effects.